# LDMol: Text-Conditioned Molecule Diffusion Model Leveraging Chemically Informative Latent Space

### v1

@@ Abstract
With the emergence of diffusion models as the frontline of generative models, many researchers have proposed molecule generation techniques using conditional diffusion models. However, due to the fundamental nature of a molecule, which carries highly entangled correlations within a small number of atoms and bonds, it becomes difficult for a model to connect raw data with the conditions when the conditions become more complex as natural language. To address this, here we present a novel latent diffusion model dubbed LDMol, which enables a natural text-conditioned molecule generation. Specifically, LDMol is composed of three building blocks: a molecule encoder that produces a chemically informative feature space, a natural language-conditioned latent diffusion model using a Diffusion Transformer (DiT), and an autoregressive decoder for molecule re. In particular, recognizing that multiple SMILES notations can represent the same molecule, we employ a contrastive learning strategy to extract the chemical informative feature space. LDMol not only beats the existing baselines on the text-to-molecule generation benchmark but is also capable of zero-shot inference with unseen scenarios. Furthermore, we show that LDMol can be applied to downstream tasks such as molecule-to-text retrieval and text-driven molecule editing, demonstrating its versatility as a diffusion model.
@@ 1 Introduction
Designing compounds with the desired characteristics is the essence of solving many chemical tasks. Inspired by the rapid development of generative models in the last decades, de novo molecule generation via deep learning models has been extensively studied. Diverse models have been proposed for generating molecules that agree with a given condition on various data modalities, including string representations [44], molecular graphs [28], and point clouds [20]. The attributes controlled by these models evolved from simple chemical properties [33, 12] to complex preferences such as biological activity [48] and multi-objective conditioning [26, 5]. More recently, as AI models’ natural language comprehension ability has rapidly increased, molecule generation aims to be controlled by natural language [29, 49].
@@
Meanwhile, diffusion models [46, 19] have emerged as a frontline of generative models over the past few years. Through a simple and stable training objective of predicting noise from noisy data [19], diffusion models have achieved highly realistic and controllable data generation [8, 22, 18]. Furthermore, leveraging that the score function of the data distribution is learned in their training [47], state-of-the-art image diffusion models enabled various applications on the image domain [43, 23, 7]. Inspired by the success of diffusion models, several papers suggested diffusion-based molecule generative models on various molecule domains including a molecular graph [31], strings like Simplified Molecular-Input Line-Entry System (SMILES) [13], and point cloud [20].
@@
Figure 1: Different strategies of data domain selection for molecule generative models. (a) The model directly learns the raw representation of the molecule such as Simplified Molecular-Input Line-Entry System (SMILES) tokens, point cloud coordinates, etc. (b) An autoencoder can be employed to let the generative model learn the latent distribution, but these latent spaces are generally hard to learn without any regularization on the autoencoder. (c) A regularized, chemically pre-trained encoder can provide a latent space learnable by external generative models.
@@
However, most molecule diffusion models presented so far have used a few, relatively simple conditions to control. Also, the diffusion model was mostly trained on a raw data representation (Figure 1-(a)). In any molecule modality, the molecule’s vast and complex correlations are encapsulated between relatively small dimensions compared to common data domains like images. As a result, the output of the generative model trained on raw molecule data often failed to faithfully follow the given conditions or showed poor data quality (e.g., low validity) as the condition became more sophisticated like natural language. Furthermore, raw molecule modality has inevitable discreteness (e.g., atom and bond type, connectivity, SMILES tokens) which makes it hard to connect the deeply studied concept of continuous diffusion models to molecule generation. Naively converting discrete values into real numbers wouldn’t provide rich information within those components, and still retain certain restrictions on the data.
These issues lead us to the idea of molecule generation via latent diffusion models, which have two major advantages: Building a diffusion model on a latent space allows the model to move out of the regulated raw data domain and work on in a minimally restricted continuous domain. Furthermore, by having the encoder pre-construct a chemically useful and interpretable feature space, the diffusion model can more easily connect the molecule data with the highly complicated condition of natural texts. In fact, we show that a naive autoencoder cannot provide a feature space that the diffusion model can learn (Figure 1-(b)), and making the encoder provide a chemically informative molecule feature (Figure 1-(c)) is essential to the performance of the latent diffusion model (see Section 4.3).
Therefore, we proposed a Latent Diffusion Molecule generative model (LDMol) that enables a text-conditioned molecule generation, trained on the latent space of the separately pre-trained molecule encoder. To construct a feature space that is chemically meaningful and easily learnable, we suggest a novel contrastive learning strategy of training the encoder to encode a structural similarity of the molecule. The experimental results show that LDMol can generate molecules that are valid and also meet the input text condition. Additionally, utilizing the modeled score function, LDMol has the potential to be applied to different multi-modal downstream tasks such as molecule-to-text retrieval, text-guided molecule editing, etc. We summarize the contribution of this work as follows:
@@
  • We propose a latent diffusion model LDMol for text-conditioned molecule generation to generate more valid molecules that are better aligned to the text condition. This approach demonstrates the potential of generative models for chemical entities in a latent space.
@@
  • We report the importance of preparing a chemically informative latent space for the molecule latent diffusion model, and suggest a novel contrastive learning method to train an encoder that captures the molecular structural similarity.
@@
  • LDMol outperforms the text-to-molecule generation baselines, and its modeled conditional distribution allows LDMol to retain the advanced attributes of diffusion models including various applications like molecule-to-text retrieval and text-driven molecule editing.
@@ 2 Background
Diffusion generative models. Diffusion models first define a forward process that perturbs the original data, and generates the data from the known prior distribution by the learned reverse process of the pre-defined forward process. For example, DDPM [19] fixed their forward process by gradually adding Gaussian noise to the data, which can be formalized as follows:
@@
Once $\theta$ is trained, novel data can be generated with the learned reverse process $p_{\theta}$: starting from the random noise $x_{T}\sim\mathcal{N}(0,{I})$, the output can be gradually denoised according to the modeled distribution of $p_{\theta}(x_{t-1}|x_{t})$.
Various real-world data generation tasks require to generate data $x_{0}$ with a given condition $c$. To build diffusion models that can generate data from the conditional data distribution $q(x_{0}|c)$, the model that predicts the injected noise should be also conditioned by $c$.
@@
Although the reverse diffusion process with ancestral sampling of $x_{0}\sim q(x_{0}|y)$ can be identical to unconditional diffusion models,  [18] proposed classifier-free guidance to improve the quality of the conditional diffusion models’ output. Specifically, using a scalar guidance scale $\omega$, classifier-free guidance sampling utilizes a linearly combined noise of conditional and unconditional model output,
$\hat{\epsilon_{\theta}}(x_{t},t,c)=\epsilon_{\theta}(x_{t},t,\emptyset)+\omega\{\epsilon_{\theta}(x_{t},t,c)-\epsilon_{\theta}(x_{t},t,\emptyset)\}$ (6)
where $\emptyset$ denotes the special condition input for unconditional noise.
@@
As the diffusion model’s target data domain becomes more complex and high-dimensional, diffusion models began to perform their diffusion process on the latent data space instead of the raw data [51, 39]. Specifically, using a pre-trained pair of an encoder $\mathcal{E}$ and its corresponding decoder $\mathcal{D}$, datapoint $x$ is encoded into $z:=\mathcal{E}(x)$ that is used to train the diffusion model. Generated latent $z^{\prime}$ from the pre-trained diffusion model is transformed back to $x^{\prime}:=\mathcal{D}(z^{\prime})$ on the original data domain.
Conditional molecule generation. Compared to more common domains for generative models like images or text, generative models for molecules are more challenging since there are rigid constraints on the data modality itself. String representations like SMILES has a strict syntax, which often makes SMILES-based molecule generative models show low validity on their output. A similar challenge also exists for the modalities of point clouds and graphs to ensure the model output represents a valid molecule.
@@
As a promising tool for many important chemical and engineering tasks like de novo drug discovery and material design, conditional molecule generation has been extensively studied with various models including recurrent neural networks [44], graph neural networks [28], and variational autoencoders [14, 27]. With the advent of large and scalable pre-trained models with transformers [53], the controllable conditions became more abundant and complicated [1, 5], and reached a text-guided molecule generation [10, 49, 29] which requires a deep comprehension ability for natural language.
@@
Recent works attempted to import the success of the diffusion model into molecule generation. Several graph-based and point cloud-based works have built conditional diffusion models that could generate molecules with simple physicochemical and biological properties [20, 31, 50]. The most related model to our work would be TGM-DLM [13], which attempted a text-conditioned molecule diffusion model trained on the sequence of tokenized SMILES indices.
3 Method
3.1 Chemically informative SMILES authoencoder
The primary goal of introducing autoencoders for image latent diffusion models is to map raw images into a low-dimensional space which reduces the computation cost [51, 39]. This is plausible because a high-resolution image has an enormous dimension on a pixel domain, yet each pixel contains little information.
In molecule generation, we chose the SMILES string as a base modality to represent molecules, and built a SMILES encoder to map raw SMILES strings into a latent vector. In this case, the role of our SMILES autoencoder has to be different from the autoencoders for images. A molecule structure can be fully expressed by a sequence of $L$ integers for SMILES tokens, where $L$ is the maximum token length. However, each token carries a significant amount of information, and hidden interactions between these tokens are much more complicated than interactions between image pixels. Therefore, the SMILES encoder should be more concerned with extracting chemical meaning into the latent space rather than reducing dimensions, even if it results in a latent space whose dimension is larger than the raw SMILES string. In fact, when a SMILES autoencoder was trained with naive reconstruction loss only, we observed that diffusion models failed to learn the latent space from this autoencoder (as will be seen in Section 4.3). This phenomenon also occurs even in the image latent to some extent, which is why autoencoders for image diffusion models are usually trained with regularizations such as KL divergence loss [17] or vector quantization [52].
Accordingly, here we propose three conditions that our SMILES autoencoder’s latent space has to satisfy: the input should be reconstructible, embed chemical information that diffusion models can learn, and have a small dimension as possible. A number of molecule encoders have been presented that can extract various useful chemical features [54, 29] including biochemical activity or human-annotated descriptions [55, 30]. Nonetheless, these molecule feature extractors didn’t aim to encode the entire information of the molecule structure therefore the input cannot be restored from the model output. Considering that an input molecule has to be reconstructed from the latent, the most important characteristic that our SMILES latent should have is that all SMILES representing the same molecule must be encoded into the same vector.
@@ 3.1 Chemically informative SMILES authoencoder
Figure 2: Overview of the proposed molecule autoencoder and the latent diffusion model. (a) SMILES encoder is trained with contrastive learning to construct latent space that embeds a structural similarity. (b) After the SMILES encoder is prepared, a compression linear layer and the autoregressive decoder are trained to restore the encoder input. (c) Training and inference process of the latent diffusion model conditioned by the output of the frozen external text encoder.
@@
Encoder design. In this respect, we trained our SMILES encoder with contrastive learning similar to simCLR [6](Figure 2-(a)). Contrastive learning aims to learn better representation by assimilating features containing similar information (i.e. positive pair) and distancing semantically unrelated features (i.e. negative pair). Taking advantage of the SMILES’s characteristic that a single molecule can be represented with many different SMILES strings, we defined two different SMILES from the same molecule as a positive pair. In the process of matching the outputs of two SMILES in a positive pair, the encoder learns the syntax and the information within SMILES so that its output encapsulates the structural information of the input molecule.
@@
Specifically, a SMILES string is fed into the SMILES encoder with a special "[SOS]" token which denotes the start of the sequence. For a batch of $N$ input SMILES $M=\{m_{1},m_{2},\dots,m_{k},\dots,m_{N}\}$, we prepare a positive pair SMILES $m_{k}^{\prime}$ for each $m_{k}$ by SMILES augmentation to construct $M^{\prime}=\{m_{1}^{\prime},m_{2}^{\prime},\dots,m_{k}^{\prime},\dots,m_{N}^{\prime}\}$. After $M$ and $M^{\prime}$ are passed through the SMILES encoder, we fed each SMILES’ output vector corresponding to the [SOS] token into an additional linear projection and normalization layer, denoting the output features as $v_{k}$ and $v_{k}^{\prime}(k=1,2,\dots,N)$.
Assimilating $v_{k}$ and $v_{k}^{\prime}$ from the positive pairs and distancing the others can be done by minimizing the following contrastive loss, which is essentially cross-entropy loss where the logit is the cosine similarity between given pairs.
@@
While our encoder was pre-trained to map structurally similar molecules to similar vectors, we found that the encoder tried to map stereoisomers to almost the same point in the latent space; indeed, in terms of structural differences, stereoisomers are the most challenging molecules to distinguish. To resolve this, if the training data has a stereoisomer, we considered it as ”hard-negative" samples and included it in the loss calculation batch with the original data.
Compressing the latent space. The pre-trained SMILES encoder maps a molecule into a vector of size $\{L\times d_{enc}\}$, where $d_{enc}$ is a feature size for the encoder. To avoid the curse of dimensionality and construct a more learnable feature space for diffusion models, we additionally employed a linear compression layer (Figure 2-(b)) to reduce the dimension size from $\{L\times d_{enc}\}$ to $\{L\times d_{z}\}$. The range of this linear layer output $z$ is a target domain of our latent diffusion model.
Decoder design. We prepared the SMILES decoder to reconstruct the input SMILES sequence from a given $z$. Following many major works that treated SMILES as a variant of language data[44, 5, 41], we built an autoregressive transformer [53] as our decoder (Figure 2-(b)) which is widely used to successfully generate sequential data [4]. Specifically, starting from the [SOS] token, the decoder predicts the next SMILES token using information from $z$ with cross-attention layers. When $m=\{t_{0},t_{1},\dots,t_{n}\}$ is the token sequence of the input SMILES, the decoder is trained to minimize the next-token prediction loss described as Eq. (9):
$L_{dec}=-\sum^{n}_{i=1}\log p(t_{n}|t_{0:n-1})\;$ (9)
Here, the decoder and the compression layer are jointly trained while the encoder’s parameter is frozen. After being fully trained, the decoder was able to reconstruct roughly 98% of the SMILES encoder input.

### v2

# LDMol: Text-to-Molecule Diffusion Model with Structurally Informative Latent Space

@@ Abstract
With the emergence of diffusion models as the frontline of generative models, many researchers have proposed molecule generation techniques with conditional diffusion models. However, the unavoidable discreteness of a molecule makes it difficult for a diffusion model to connect raw data with highly complex conditions like natural language. To address this, we present a novel latent diffusion model dubbed LDMol for text-conditioned molecule generation. LDMol comprises a molecule autoencoder that produces a learnable and structurally informative feature space, and a natural language-conditioned latent diffusion model. In particular, recognizing that multiple SMILES notations can represent the same molecule, we employ a contrastive learning strategy to extract feature space that is aware of the unique characteristics of the molecule structure. LDMol outperforms the existing baselines on the text-to-molecule generation benchmark, suggesting a potential for diffusion models can outperform autoregressive models in text data generation with a better choice of the latent domain. Furthermore, we show that LDMol can be applied to downstream tasks such as molecule-to-text retrieval and text-guided molecule editing, demonstrating its versatility as a diffusion model.
@@ 1 Introduction
Designing compounds with the desired characteristics is the essence of solving many chemical tasks. Inspired by the rapid development of generative models in the last decades, de novo molecule generation via deep learning models has been extensively studied. Diverse models have been proposed for generating molecules that agree with a given condition on various data modalities, including string representations (Segler et al., 2017), molecular graphs (Lim et al., 2020), and point clouds (Hoogeboom et al., 2022). The attributes controlled by these models evolved from simple chemical properties (Olivecrona et al., 2017; Gómez-Bombarelli et al., 2018) to complex biological activity (Staszak et al., 2022) and multi-objective conditioning (Li et al., 2018; Chang & Ye, 2024). More recently, as deep learning models’ natural language comprehension ability has rapidly increased, there’s a growing interest in molecule generation controlled by natural language (Edwards et al., 2022; Pei et al., 2023; Liu et al., 2024a; Su et al., 2022) which encompasses much broader and user-friendly controllable conditions.
@@
Meanwhile, diffusion models (Song & Ermon, 2019; Ho et al., 2020) have emerged as a frontline of generative models over the past few years. Through a simple and stable training objective of predicting noise from noisy data (Ho et al., 2020), diffusion models have achieved highly realistic and controllable data generation (Dhariwal & Nichol, 2021; Karras et al., 2022; Ho & Salimans, 2021). Furthermore, leveraging that the score function of the data distribution is learned in their training (Song et al., 2021b), state-of-the-art image diffusion models enabled various applications on the image domain (Saharia et al., 2022; Kim & Ye, 2021; Chung et al., 2023). Inspired by the success of diffusion models, several papers suggested diffusion-based molecule generative models on various molecule domains including a molecular graph (Luo et al., 2023), strings like Simplified Molecular-Input Line-Entry System (SMILES) (Gong et al., 2024), and point clouds (Hoogeboom et al., 2022).
@@
Figure 1: Different strategies of data domain selection for molecule diffusion models. (a) The model directly learns the raw representation of the molecule such as string format, point cloud coordinates, etc. (b) An autoencoder can be employed to let the generative model learn its latent distribution. (c) A regularized, chemically pre-trained encoder can provide a latent space learnable by external generative models.
@@
However, a discrepancy between molecule data and common data domains like images makes it hard to connect the diffusion models to molecule generation. Whereas diffusion models are deeply studied on a continuous data domain with gaussian noise, any molecule modality has inevitable discreteness such as atom and bond type, connectivity, and SMILES tokens as shown in Figure 1-(a). Naively converting discrete values into real numbers would ignore rich correlation within the components, and still retain certain restrictions on the data. As a result, diffusion models trained on raw molecule data often failed to faithfully follow the given conditions or showed poor data quality (e.g., low validity) as the condition became more sophisticated like natural language. Most molecule diffusion models presented so far have used a few, relatively simple conditions to control, while major developments in text-conditioned molecule generative models were based on autoregressive models.
To overcome the limitation of the previous works (Xu et al., 2023) that mainly focused on resolving the discreteness of the molecular domain space by training with reconstruction loss (Figure 1-(b)), we aim to build a latent space encoder to extract rich and refined information about the molecule structure such as Figure 1-(c). Specifically, we design a novel Latent Diffusion Molecular generative model (LDMol) for text-conditioned molecule generation, trained on the latent space of the separately pre-trained molecule encoder. By preparing an encoder to provide a chemically useful and interpretable feature space, our model can more easily connect the molecule data with the highly complicated condition of natural texts. In the process, we suggest a novel contrastive learning strategy of training a molecular encoder to encode a structural invariant of the molecule.
The experimental results show that LDMol can outperform many state-of-the-art autoregressive models and generate valid SMILES that meet the input text condition. Considering SMILES as a variation of text data, we report one of the first diffusion models that successfully surpassed autoregressive models in text data generation. This suggests the possibility of improving existing diffusion models (Lovelace et al., 2024) for natural language through careful design of the latent space. Furthermore, LDMol can leverage the learned score function and be applied to several multi-modal downstream tasks such as molecule-to-text retrieval and text-guided molecule editing, without additional task-specific training. We summarize the contribution of this work as follows:
@@
  • We propose a latent diffusion model LDMol for text-conditioned molecule generation to generate valid molecules that are better aligned to the text condition. This approach demonstrates the potential of generative models for chemical entities in a latent space.
@@
  • We report the importance of preparing a chemically informative latent space for the molecule latent diffusion model, and suggest a novel contrastive learning method to train an encoder that captures the molecular structural characteristic.
@@
  • LDMol outperforms the text-to-molecule generation baselines, and its modeled conditional score function allows LDMol to retain the advanced attributes of diffusion models including various applications like molecule-to-text retrieval and text-guided molecule editing.
@@ 2 Background
Diffusion generative models. Diffusion models first define a forward process that perturbs the original data, and generates the data from the known prior distribution by the learned reverse process of the pre-defined forward process. Ho et al. (2020) fixed their forward process by gradually adding Gaussian noise to the data, which can be formalized as follows:
@@
Once $\theta$ is trained, novel data can be generated with the learned reverse process $p_{\theta}$; starting from the random noise $x_{T}\sim\mathcal{N}(0,{I})$, the output can be gradually denoised according to the modeled distribution of $p_{\theta}(x_{t-1}|x_{t})$.
Various real-world data generation tasks require to generate data $x_{0}$ with a given condition $c$. To build diffusion models that can generate data from the conditional data distribution $q(x_{0}|c)$, the model that predicts the injected noise should also be conditioned by $c$.
@@
Diffusion models can perform their diffusion process on the latent data space instead of the raw data when the raw data is too complex or high-dimensional (Vahdat et al., 2021; Rombach et al., 2022). Specifically, using a pre-trained pair of an encoder $\mathcal{E}(\cdot)$ and its corresponding decoder $\mathcal{D}(\cdot)$, datapoint $x$ is encoded into $z:=\mathcal{E}(x)$ that is used to train the diffusion model. Generated latent $z^{\prime}$ from the pre-trained diffusion model is transformed back to $x^{\prime}:=\mathcal{D}(z^{\prime})$ on the original data domain. The latent diffusion model was also adopted for discrete texts (Lovelace et al., 2024), where the latent space was constructed by a denoising autoencoder.
@@
Conditional molecule generation. As a promising tool for many important chemical and engineering tasks like de novo drug discovery and material design, conditional molecule generation has been extensively studied with various models including recurrent neural networks (Segler et al., 2017), graph neural networks (Lim et al., 2020), and variational autoencoders (Gómez-Bombarelli et al., 2018; Lim et al., 2018). With the advent of large and scalable pre-trained models with transformers (Vaswani et al., 2017), the controllable conditions became more abundant and complicated (Bagal et al., 2021; Chang & Ye, 2024), and reached a text-guided molecule generation (Edwards et al., 2022; Su et al., 2022; Liu et al., 2024a) which requires a deep comprehension ability for natural language.
@@
Recent works attempted to import the success of the diffusion model into molecule generation. Several graph-based and point cloud-based works have built conditional diffusion models that could generate molecules with simple chemical and biological conditions (Hoogeboom et al., 2022; Luo et al., 2023; Trippe et al., 2023). Gong et al. (2024) attempted a text-conditioned molecule diffusion model trained on the sequence of tokenized SMILES indices. However, these models treated discrete molecule attributes with continuous Gaussian diffusion models, introducing arbitrary numeric values and suboptimal performances. Xu et al. (2023) employed an autoencoder to build a diffusion model on a smooth latent space, but its controllable conditions were still limited to several physiochemical properties.
3 Methods
In this section, we explain the overall model architecture and training procedure of the proposed LDMol, which are briefly illustrated in Figure 2. We first discuss the conditions the molecule encoder should meet and present our strategy to construct an encoder that can extract structural information from the raw molecule representation. Then we describe how we prepared a complete encoder-decoder pair with this encoder and trained a text-conditioned diffusion model on its latent space.
@@ 3 Methods
Figure 2: Overview of the proposed molecule autoencoder and the latent diffusion model. (a) SMILES encoder is trained with contrastive learning to construct latent space that embeds a structural characteristic. (b) After the SMILES encoder is prepared, a linear compression layer and an autoregressive decoder are trained to restore the encoder input. (c) The training and inference process of the latent diffusion model is conditioned by the output of the frozen external text encoder.
3.1 Chemically informative SMILES autoencoder
The primary goal of introducing autoencoders for image latent diffusion models is to map raw images into a low-dimensional space which reduces the computation cost (Vahdat et al., 2021; Rombach et al., 2022). This is plausible because a high-resolution image has an enormous dimension on a pixel domain, yet each pixel contains little information.
In this work for molecule generation, we utilized SMILES, one of the most widely used molecule representations in chemical databases and benchmarks, and built a SMILES encoder to map raw SMILES strings into a latent vector. In this case, the role of our SMILES autoencoder has to be different from that of the autoencoders for images; a molecule structure can be fully expressed by only a sequence of $L$ integers for SMILES tokens, where $L$ is the maximum token length. However, each token carries significant information, and hidden interactions between these tokens are much more complicated than interactions between image pixels. Therefore, the SMILES encoder should focus more on extracting chemical meaning into the latent space, even if it results in a latent space with more dimension than the raw SMILES string.
A number of molecule encoders (Wang et al., 2019; Liu et al., 2024a; Zeng et al., 2022; Liu et al., 2023a) have been presented that can extract various useful chemical features including biochemical activity or human-annotated descriptions. Nonetheless, these molecule encoders aim to extract certain desired features rather than encode all the information about the molecule structure. Therefore the input cannot be fully restored from the model output.
Although autoencoders with appropriate regularization (e.g., KL-divergence loss (Kingma, 2013; Gómez-Bombarelli et al., 2018)) provide a continuous and reconstructible molecular latent space, their encoder output is not guaranteed to possess the characteristic of the underlying molecular structure, beyond the minimal information to reconstruct the input string. To visualize this, we prepared a trained $\beta$-VAE (Higgins et al., 2017) and measured the feature distance between two SMILES from the same molecule obtained via SMILES enumeration (Bjerrum, 2017). Here, SMILES enumeration is the process of writing out all possible SMILES of the same molecule, as illustrated in Figure  3-(a). Figure  3-(b) shows that $\beta$-VAE had difficulties assimilating features from the same molecule compared to the one between random SMILES pairs, indicating that it couldn’t capture the intrinsic features beneath the SMILES string. This inherited inconsistency makes it difficult for later models that learn this latent space to figure out the connection between the latent and the molecule, which could eventually degrade the performance of the generative model with complex conditions. If a molecule encoder can truly provide a unique representation of the atoms and their connectivity, all SMILES representing the same molecule must be encoded into the same output and should be differentiated from the other random SMILES.
Figure 3: Behaviour of encoder features on SMILES enumeration. (a) Examples of SMILES enumeration with node traversal order. (b) Euclidean distance between features from $\beta$-VAE ($\beta=0.001$) and our proposed encoder, with 1,000 random SMILES pairs and 1,000 enumerated SMILES pairs. The distance was rescaled by $1/\sqrt{d}$ where $d$ is a latent dimension size.
Accordingly, here we propose three conditions that our SMILES autoencoder’s latent space has to satisfy: enable reconstruction of the input, have as small dimensions as possible, and embed chemical information that can be readily learned by diffusion models.
@@ 3.1 Chemically informative SMILES autoencoder
Encoder design. In this respect, we trained our SMILES encoder with contrastive learning (Figure 2-(a)), which aims to learn better representation by assimilating features containing similar information (i.e. positive pair) and distancing semantically unrelated features (i.e. negative pair). We defined two enumerated SMILES from the same molecule as a positive pair and two SMILES from different molecules as a negative pair.
Here, we argue that the proposed contrastive learning with SMILES enumeration can train the encoder to encapsulate the structural information of the input molecule. Contrastive learning learns invariant for the augmentations applied on positive pairs (Zhang & Ma, 2022), and it is known that a good augmentation should reduce as much mutual information between positive pairs as possible while preserving relevant information (Tian et al., 2020). Meanwhile, enumerated SMILES of the same molecule are obtained by traversing the nodes and edges in the molecular graph with a different visiting order. Therefore, to detect all possible enumerated SMILES and find SMILES-enumeration-invariant, the model has to understand the entire connectivity between atoms. This makes the encoder output a unique characteristic that captures the overall molecule structure. Compared to the hand-crafted augmentations previously presented for molecule contrastive learning (You et al., 2020), enumerated SMILES pairs have less mutual information since we utilize all possible variations in the SMILES format. And since all enumerated SMILES are guaranteed to represent an identical molecule, there is no relevant information loss during the augmentation. Figure  3-(a) confirms that our LDMol trained with the contrastive learning with SMILES enumeration now correctly assimilates features from the same molecule compared to the one between random SMILES pairs.
@@
Specifically, a SMILES string is fed into the SMILES encoder $\mathcal{E}(\cdot)$ with a special “[SOS]” token which denotes the start of the sequence. For a batch of $N$ input SMILES $M=\{m_{1},m_{2},\dots,m_{k},\dots,m_{N}\}$, we prepare a positive pair SMILES $m_{k}^{\prime}$ for each $m_{k}$ by SMILES enumeration to construct $M^{\prime}=\{m_{1}^{\prime},m_{2}^{\prime},\dots,m_{k}^{\prime},\dots,m_{N}^{\prime}\}$. After $M$ and $M^{\prime}$ are passed through the SMILES encoder, we fed each SMILES’ output vector corresponding to the [SOS] token into an additional linear projection and normalization layer, denoting its output as $v_{k}$ and $v_{k}^{\prime}(k=1,2,\dots,N)$. Assimilating $v_{k}$ and $v_{k}^{\prime}$ from the positive pairs and distancing the others can be done by minimizing the following InfoNCE loss (Chen et al., 2020).
@@
While our encoder was pre-trained to map structurally similar molecules to similar vectors, we found that the encoder tried to map stereoisomers to almost the same point in the latent space; indeed, in terms of structural differences, stereoisomers are the most challenging molecules to distinguish. To resolve this, if the training data has a stereoisomer, we consider it as “hard-negative” samples and include it in the loss calculation batch with the original data.
Compressing the latent space. The pre-trained SMILES encoder maps a molecule into a vector of size $[L\times d_{enc}]$, where $d_{enc}$ is a feature size of the encoder. To avoid the curse of dimensionality and construct a more learnable feature space for diffusion models, we additionally employed a linear compression layer $f(\cdot)$ (Figure 2-(b)) to reduce the dimension from $[L\times d_{enc}]$ to $[L\times d_{z}]$. The range of this linear layer output is a target domain of our latent diffusion model.
Decoder design. When a SMILES $m$ is passed through the SMILES decoder and the compression layer, the SMILES decoder reconstructs $m$ from $f(\mathcal{E}(m))$. Following many major works that treated SMILES as a variant of language data (Segler et al., 2017; Chang & Ye, 2024; Ross et al., 2022), we built an autoregressive transformer (Vaswani et al., 2017) as our decoder (Figure 2-(b)) which is widely used to successfully generate sequential data (Brown et al., 2020). Specifically, starting from the [SOS] token, the decoder predicts the next SMILES token using information from $f(\mathcal{E}(m))$ with cross-attention layers. When $\{t_{0},t_{1},\dots,t_{n}\}$ is the token sequence of $m$, the decoder is trained to minimize the next-token prediction loss described as Eq. (8). Here, the decoder and the compression layer are jointly trained while the encoder’s parameter is frozen.
$\mathcal{L}_{dec}=-\sum^{n}_{i=1}\log p(t_{n}|t_{0:n-1},f(\mathcal{E}(m)))\;$ (8)
As long as the encoder output contains complete information of the molecule structure, the decoder can reconstruct the frozen encoder input. After being fully trained, the decoder was able to reconstruct roughly 98% of the SMILES encoder input.